OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 3 times x + 3 the square root of 3 times x − 2 the square root of 3 times x − 10 −10 2 times x times the square root of 3 times x 2 times the square root of 3 times x − 10 10 − 2 times x times the square root of 3 times x

OpenStudy (tkhunny):

Not clear at all what is wanted. Can you communicate it better? You have six expressions or values. Just exactly how are they related? Look down below at the \(\left[\sum\;Equation\right]\) button. You may be able to write it exactly as it appears in the question.

OpenStudy (anonymous):

\[\sqrt{3x} + 3\sqrt{3x} - 2\sqrt{3x} - 10\]

OpenStudy (tkhunny):

Could you simplify it is it said. \(Frog + 3Frog - 2Frog\)?? How many frogs?

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

It's a simple enough question. If you have a frog, you then add three frogs, and you then subtract 2 frogs, how many frogs have you?

OpenStudy (anonymous):

2.

OpenStudy (tkhunny):

Perfect. Whatever the units, be they frogs or \(\sqrt{3x}\), it is of no consequence. So, is your answer to the question \(2\sqrt{3x} - 10\)? Maybe, \(2(\sqrt{3x} - 5)\)